OpenStudy (anonymous):

Is the series convergent or divergent using comparison test: I'll type it below.

OpenStudy (anonymous):

\[\sum_{n = 1}^{\infty}e ^{1/n}/n\]

OpenStudy (anonymous):

Limit comparison or direct?

OpenStudy (anonymous):

I cant think of any thing to compare it to directly.

OpenStudy (anonymous):

Should I compare it to 1/n ?

OpenStudy (anonymous):

Well, you know that your series is always bigger than 1/n just slightly. And 1/n is the divergent harmonic so I believe thats all you need. Is that: \[\sum_{n=1}^{\infty}\frac{e^{\frac{1}{n}}}{n}>\sum_{n=1}^{\infty}\frac{1}{n}\] And 1/n diverges so your series diverges by direct comparison.

OpenStudy (anonymous):

Ha, it is really that simple. Thank you!

OpenStudy (anonymous):

No problem :P